OpenStudy (anonymous):

Solve the following system of equations. Explain which method you used and why you chose that method 10x-6y=2 5x+3y=5

OpenStudy (mathstudent55):

I have a suggestion. The addition method is practical to use when, by adding the equations, a variable is eliminated.

OpenStudy (mathstudent55):

That can be done easily here. Notice the first equation has the term -6y and the second equation has the term 3y. If you turn the term 3y of the second equation into 6y, then adding those terms would eliminate y, since 6y + (-6y) = 0. Since you need 6y in the second equation, multiply the entire second equation by 2. Then add the equations together.

OpenStudy (mathstudent55):

Copy the first equation exactly as it is. Below it, write the second equation multiplied by 2. Then add the two equations.

OpenStudy (anonymous):

what about 2 and 5, do I add both of them together too?

OpenStudy (mathstudent55):

Yes, but remember that the 5 becomes 10 when you multiply the second equation by 2. I'll show it to you below. Here is the original system of equations: \(10x-6y=2 \) \(5x+3y=5\) Now we copy the first equation as it is. Below it, we write the second equation after multiplying it by 2 on both sides. \(10x-6y=2 \) \(10x+6y=10\) Now we add the two equations: \(20x + 0y = 12\) \(20x = 12\) Now that we have one equation in one variable, x, we solve for the variable. Divide both sides by 20: \(\dfrac{20}{20}x = \dfrac{12}{20}\) Reduce the fraction by dividing the numerator and denominator by 5. \(x = \dfrac{3}{5} \) Now that we have x, we can use the equations to eliminate x and solve for y. Let's go back to the step where the second equation had been multiplied by 2. The two equations are (copied form above): \(10x-6y=2 \) \(10x+6y=10\) Now we subtract the equations. The terms 10x in both equations will be eliminated leaving an equation in only y. Subtracting the equations we get: \(-12y = -8\) Divide both sides by -12 and reduce the fration: \(\dfrac{-12}{-12}y = \dfrac{-8}{-12}\) \(y = \dfrac{8}{12} \) \(y = \dfrac{2}{3} \) The solution is: \(x = \dfrac{3}{5} \), \(y = \dfrac{2}{3} \)